OpenStudy (anonymous):

The speed of sound V near the Earth's surface can be represented by this equation, V=20 SQRT t + 273 , where t is the temperature in degrees Celsius. Find the temperature when V = 380 meters per second. Round to the nearest whole number. A. 88 B. 90 C. 92 D. 94

OpenStudy (anonymous):

pls help!! :P

OpenStudy (anonymous):

plug in v, divide by 20, then square

OpenStudy (anonymous):

well plug in the value of V and simply solve for t :)

OpenStudy (anonymous):

Yep. \[V = 380 \implies 380 = 20\sqrt{t+237}\]

OpenStudy (anonymous):

ohhhh so its just like all the other problems i have been posting wow lol i hate story problems period tho:P

OpenStudy (anonymous):

380=20 SQRT t + 273 107=20sqrt t (107/20)^2=t

OpenStudy (anonymous):

Usually each sentence will be giving you an equation or inequality that you can use. Just try to interpret the sentence into an equation, then take all the equations and try to find what they're looking for.

OpenStudy (anonymous):

ahhhh alright that helps a bunch gave all of ya guys medal thingys!! (: